I need help solving I'm not sure where to start.
The question is a retangular pool is 4ft. longer than it is wide. A small concrete surrounds the pool. The walkway is a constant 1ft and has an area of 72ft^2. What are the dimensions of the pool?

Answer by stanbon(75887)

The question is a retangular pool is 4ft. longer than it is wide. A small concrete surrounds the pool. The walkway is a constant 1ft and has an area of 72ft^2. What are the dimensions of the pool?
-----------------------------------------
Draw a rectangle inside a rectangle.
The inner rectangle has dimentions x and x+4 and area = x^2+4x sq. ft
------
The outer rectangle has dimentions x+2 and x+6 and area = x^2+8x+12 sq ft.
======
Equations
Larger area - Smaller area = 72 sq ft
------
(x^2+8x+12)-(x^2+4x) = 72 sq ft
4x+12 = 72
4x = 60
x = 15 ft (width of the pool)
x+4 = 19 ft (length of the pool)
==================================
Cheers,
Stan H.
----